Evaluate the discriminant of each equation.How many real and imaginary solutions does it have.
1.4x^2-4x+1=0
2.x^2-6x+7=0

Answer by stanbon(75887)

Evaluate the discriminant of each equation.How many real and imaginary solutions does it have.
1.4x^2-4x+1=0
b^2-4ac = 16-4*4*1 = 0
So, the equation has 2 equal Real Number solutions.
---------------------------------
2.x^2-6x+7=0
b^2-4ac = 36-4*1*7 = 8
Since the determinant is positive there are
2 unequal Real Number solutions.
===================================
Cheers,
Stan H.